# Configuring Agentic AI Coding Tools: An Exploratory Study

### v1

@@ Abstract.
Agentic AI coding tools with autonomous capabilities beyond conversational content generation increasingly automate repetitive and time-consuming software development tasks. Developers can configure these tools through versioned repository-level artifacts such as Markdown and JSON files. In this paper, we present a systematic analysis of configuration mechanisms for agentic AI coding tools, covering Claude Code, GitHub Copilot, Cursor, Gemini, and Codex. We identify eight configuration mechanisms and, in an empirical study of 2,926 GitHub repositories, examine whether and how they are adopted. We then explore Context Files, Skills, and Subagents, that is, three mechanisms available across tools, in more detail. Our findings reveal three trends. First, Context Files dominate the configuration landscape and are often the sole mechanism in a repository, with AGENTS.md emerging as an interoperable standard across tools. Second, advanced mechanisms such as Skills and Subagents are only shallowly adopted: most repositories define only one or two artifacts, and Skills predominantly rely on static instructions rather than executable workflows. Third, distinct configuration cultures are forming around different tools, with Claude Code users employing the broadest range of mechanisms. These findings establish an empirical baseline for longitudinal and experimental research on how configuration strategies evolve and affect agent performance as agentic AI coding tools mature.
@@ 1. Introduction
Agentic AI coding tools based on large language models (LLMs) (Sergeyuk et al., 2025; Terragni et al., 2025), automate time-consuming and repetitive tasks, such as generating and editing code, tests, and documentation. The agentic capabilities of these tools go beyond those of purely reactive conversational assistants. Instead of just following user commands, these tools take initiative and act proactively to accomplish defined objectives (Anthropic, 2025). They can autonomously interact with development environments and use project artifacts, external data, and command-line tools to perform actions with reduced human interaction (Yang et al., 2024). Certain functionality can be delegated to “tools” and “(sub)agents.” Tools are capabilities with a specific bounded function (e.g., run_python(code) or search_web(query)). [^1] We distinguish such “tools” from full-fledged “agentic tools” (e.g., Claude Code). They are deterministic and invoked by the underlying model or agents. (Sub)agents are goal-directed action-taking components. They interpret a user goal, break it down into substeps, choose appropriate tools, execute actions, observe results, and potentially adjust their plan for the next iteration of this “agent loop.” Initially, agentic AI coding tools implemented one central agent loop steered by a foundational model such as Anthropic’s Opus or OpenAI’ GPT. Example tools include Claude Code and OpenAI Codex. Soon, conversational tools such as GitHub Copilot and Cursor offered similar capabilities in an “agent mode.” Over the months preceding the writing of this paper, tool vendors implemented extension and configuration mechanisms that allow developers, for example, to write their own (sub)agents that operate in parallel to the central agent loop.
We define configuration mechanisms as means by which developers can tailor tool and agent behavior to a project or workflow (e.g., by providing contextual information in a file or by specifying dedicated subagents). A configuration artifact is the tangible specification of that mechanism (e.g., the actual Markdown file with the contextual information). For example, files like AGENTS.md or CLAUDE.md act as “READMEs for agents” with context-specific information about build commands, coding conventions, and rules for CI/CD pipelines (Mohsenimofidi et al., 2026; OpenAI, 2025). Configuration artifacts are version-controlled, inspectable, and collaboratively maintained. With the increased availability and diversity of agentic tools, the number of configuration mechanisms and related artifacts has also increased.
@@
With this paper, we provide the following contributions: (1) We systematically documented eight configuration mechanisms: Context Files, Skills, Subagents, Commands, Rules, Settings, Hooks, and MCP servers. We identified these from the documentation of Claude Code, GitHub Copilot, Cursor CLI, Gemini CLI and Codex CLI. Some mechanisms are available in all tools; others are specific to particular tools. (2) We analyzed the adoption of these configuration mechanisms in 2,926 GitHub repositories. Context Files (Markdown files that provide contextual project information) dominated and are often the sole configuration mechanism. Claude Code users apply the broadest range of configuration mechanisms. (3) We analyzed the adoption of Context Files, Skills (bundles of information and tools), and Subagents (specialized agents with their own model context), in more detail. Three Context File formats showed the most dynamic development, with AGENTS.md emerging as a standard. Although Skills and Subagents can be defined for a wide range of purposes, most repositories that adopt them define only one or two artifacts. Moreover, Skills primarily rely on static resources rather than executable scripts to extend tool and agent behaviour.
@@ 2. Related Work
Related work has examined how repository-level context files, prompts, and structured context mechanisms shape agent behavior, efficiency, and integration into development workflows. Although several studies have investigated individual configuration artifacts, most notably repository-level context files such as AGENTS.md, prior work has typically focused on single artifact types in isolation. There has been no systematic examination of the broader configuration mechanisms to customize agentic behavior. Independent of tool architecture, context engineering allows developers to design, structure, and supply task- and project-specific information to guide the behavior of tools (Mei et al., 2025; Schmid, 2025; Horthy, 2025; Sahoo et al., 2024; DAIR.AI Prompt Engineering Guide, 2025). Empirical work has mainly focused on repository-maintained context files, such as AGENTS.md. Mohsenimofidi et al. (Mohsenimofidi et al., 2026) analyzed the emergence of context engineering practices in open-source projects and showed that repository-level instruction files serve as persistent configuration mechanisms that encode architectural constraints, build commands, and workflow conventions. Complementing this work, Chatlatanagulchai et al. (Chatlatanagulchai et al., 2025) found that context files are actively maintained, structurally consistent, and focus predominantly on functional development instructions such as build and testing procedures. Santos et al. (Santos et al., 2025) analyzed Claude Code projects and how configuration artifacts structure agent interactions in repositories.
@@
In other work, Villamizar et al. (Villamizar et al., 2025) argued that prompts should be treated as software engineering artifacts and outlined a research agenda to understand their evolution, reuse, and governance. Beyond static prompts, Zhang et al. (Zhang et al., 2025) introduced a framework that treats contexts as evolving instructions and addresses issues such as brevity bias and context collapse through incremental updates.
Recent work further generalized this perspective by formalizing context as a programmable abstraction. Zhang and Wang (Zhang and Wang, 2025) introduced Monadic Context Engineering, modeling context construction and transformation through monadic composition. McMillan (McMillan, 2026) evaluated structured file-native context schemas and showed how format and multi-file organization influence retrieval accuracy and downstream agent performance. Ye et al. (Ye et al., 2026) proposed Meta Context Engineering, a framework in which agentic skills and context artifacts co-evolve via evolutionary search.
However, a cross-tool and cross-agent perspective mapping configuration mechanisms and studying their adoption patterns across repositories is missing. Our work addresses this gap.
@@ 3. Data Collection and Analysis
We collected OSS projects hosted on GitHub. Since GitHub hosts a variety of content, we selected repositories belonging to “engineered” software projects (Munaiah et al., 2017). We adopted and updated a selection approach used in prior work, starting with the SEART GitHub search tool  (Mohsenimofidi et al., 2026; SEART, 2025; Dabic et al., 2021). We selected non-fork repositories with at least two contributors and a license that were created before 1 January 2024 with commits since 1 June 2025. We then excluded archived, disabled, or locked repositories We applied a licensing filter and a language filter based on popularity, selecting Python, TypeScript, JavaScript, Go, Java, C++, Rust, PHP, C#, and C, and applied filters based on commit and watcher count based on prior research (Mohsenimofidi et al., 2026). This resulted in a final sample of 37,249 repositories.
@@
a Copilot also supports CLAUDE.md, AGENTS.md, and GEMINI.md; b Configured via the GitHub web interface, not via files in the project repository;
@@
Table 2. Release dates of agentic AI coding tools and repositories ($n=2,926$; one repository can use multiple tools).
@@
Claude Code 02/2025 (CLI & agents since release) 1,310
@@
958
Codex CLI 04/2025 (CLI & agents since release) 558
@@
186
Table 3. Repository metadata comparison by agentic tool. Cells show median, IQR, and Cliff’s $\delta$ for tool-adopting repositories. Significant differences (BH-adjusted) are in bold.
@@
All 6.7 42 2,106 40k
(nn=2,926) (4.3–9.4) (19–105) (965–5,064) (9,498–132k)
Claude 6.2∗∗∗ 40 2,234 42k
(nn=1,310) (4.1–9.2) (19–111) (995–5,268) (10k–151k)
$\delta$=-0.08 $\delta$=0.00 $\delta$=0.03 $\delta$=0.04
Gemini 6.8 66∗∗∗ 3,322∗∗∗ 58k∗∗
(nn=186) (4.7–9.5) (32–144) (1,290–9,776) (15k–193k)
$\delta$=0.03 $\delta$=0.21 $\delta$=0.21 $\delta$=0.12
Codex 6.9 43 1,809∗∗ 34k∗∗
(nn=568) (4.2–9.3) (19–89) (877–4,058) (8,696–94k)
$\delta$=0.01 $\delta$=-0.02 $\delta$=-0.10 $\delta$=-0.08
Copilot 7.1∗∗∗ 42 2,200 44k
(nn=958) (5.0–9.7) (19–113) (1,000–5,591) (9,858–143k)
$\delta$=0.11 $\delta$=0.02 $\delta$=0.04 $\delta$=0.03
Cursor 5.5∗∗∗ 54∗∗ 2,780∗∗∗ 75k∗∗∗
(nn=355) (3.4–7.9) (23–120) (1,264–5,968) (21k–201k)
$\delta$=-0.19 $\delta$=0.10 $\delta$=0.14 $\delta$=0.22
Note: Mann–Whitney U test (two-sided) with Benjamini–Hochberg FDR correction (20 comparisons). Effect sizes: Cliff’s delta. $p<.05$; ∗∗ $p<.01$; ∗∗∗ $p<.001$ (adjusted).
Figure 2. Adoption of agentic tools per programming language; % relative to #repositories using a certain tool.
Figure 3. Usage of configuration mechanisms across agentic tools. % relative to #repositories using a certain mechanism.
@@
Figure 1 outlines our data collection pipeline for these repositories. We cloned them and searched their default branch for a README file. We excluded 11 repositories that became unavailable in the meantime, 60 repositories without a README file, and then used the lingua-language-detector Python library to detect the language of files we found. This resulted in 994 repositories with non-English README files being excluded.
For the remaining 36,184 repositories, we developed a classification pipeline that uses the GPT-5.2 model to determine, based on the content of a README file, whether the software project fulfills our definition of an “engineered” project. We based our definition on Munaiah et al.’s paper (Munaiah et al., 2017). An engineered project is a project that shows clear evidence of software engineering practices (e.g., it includes a clearly stated purpose of a problem that the software solves; it provides installation and deployment instructions, developer and user documentation, testing, CI/CD, or quality-assurance-related information). There should also be evidence for maintenance. This definition was part of our prompt, together with the README of each repository. In our supplementary material, we share the final prompt and previous iterations of the prompt, which we tested with subsets of the data, as well as the exact configuration for labeling repositories as “engineered” or not (we kept OpenAI’s default values for all parameters) (Authors, 2026). The classification pipeline labeled 32,564 repositories as engineered software projects, was ‘unsure’ about 2,204 repositories, and excluded 1,416 repositories. For each label category, we spot-checked randomly selected samples, both during the iterative development of the prompts and for the final sample.
We then cloned the remaining 32,564 repositories and applied heuristics based on file names and file paths to detect the usage of AI coding tools and configuration mechanisms (see Table 1). These heuristics are based on the answers to RQ1 and are briefly discussed in Section 4. This resulted in 2,926 repositories that use one or more AI coding tools. Our data collection and analysis scripts and the analyzed data are available online (Authors, 2026).
@@ 4. Configuration Mechanisms (RQ1)
To determine the set of configuration mechanisms for agentic AI coding tools, we followed two steps. We first selected five agentic AI coding tools based on the 2025 Stack Overflow Developer Survey (Stack Exchange Inc., 2026). The tools represent the four most popular AI tools among the surveyed developers. We added Codex instead of ChatGPT as Codex is an agentic tool by the same vendor using the same models. We added Cursor because it has recently been studied in software engineering  (He et al., 2026; Jiang and Nam, 2026). Table 2 lists the five tools. Our study focuses on agentic tools with a command-line interface (CLI) that first appeared in 2025. For some tools, prior non-agentic versions exist.
In the second step, one author systematically reviewed the online documentation of each tool, documenting configuration mechanisms together with repository-level files and directories that might indicate their usage. This was then cross-checked by two other authors and discussed among all.
As part of RQ1 and to “operationalize” configuration mechanisms and their occurrence in repositories, we then used that information to develop heuristics to identify whether tools are used in a repository (see Table 1). Note that for GitHub Copilot, Cursor, and Gemini, files and directories indicating their use apply both to their conversational and agentic interfaces. For example, a .claude folder or a CLAUDE.md file indicate usage of Claude Code. In addition, we developed strategies to determine whether configuration mechanisms were used. For example, if a directory .claude/agents/ with Markdown files is present, this indicates usage of Subagents. All documents and the Python scripts that implement our matching strategy are part of the supplementary material (Authors, 2026).
@@
Key insights for RQ1:
  • Qualitative analysis of the documentation of agentic tools revealed eight configuration mechanisms.
  • There is convergence for some mechanisms across tools (e.g., all five support Context Files and Skills), yet no single tool implements all eight mechanisms, suggesting the ecosystem is still maturing, standards are developing.

### v2

@@ Abstract.
Agentic AI coding tools increasingly automate software development tasks. Developers can configure these tools through versioned repository-level artifacts such as Markdown and JSON files. We present a systematic analysis of configuration mechanisms for agentic AI coding tools, covering Claude Code, GitHub Copilot, Cursor, Gemini, and Codex. We identify eight configuration mechanisms spanning a spectrum from static context to executable and external integrations, and, in an empirical study of 2,923 GitHub repositories, examine whether and how they are adopted, with a detailed analysis of Context Files, Skills, and Subagents. First, Context Files dominate the configuration landscape and are often the sole mechanism in a repository, with AGENTS.md emerging as an interoperable standard across tools. Second, advanced mechanisms such as Skills and Subagents are only shallowly adopted. Most repositories define only one or two artifacts, and Skills predominantly rely on static instructions rather than executable workflows. Third, distinct configuration cultures are forming around different tools, with Claude Code users employing the broadest range of mechanisms. These findings establish an empirical baseline for understanding how developers configure agentic tools, suggest that AGENTS.md serves as a natural starting point, and motivate longitudinal and experimental research on how configuration strategies evolve and affect agent performance.
@@ 1. Introduction
Agentic AI coding tools based on large language models (LLMs) (Sergeyuk et al., 2025; Terragni et al., 2025), automate time-consuming and repetitive tasks, such as generating and editing code, tests, and documentation. Unlike purely reactive conversational assistants, these tools proactively accomplish defined objectives (Anthropic, 2025) by autonomously interacting with development environments, project artifacts, external data, and command-line tools with reduced human interaction (Yang et al., 2024). Certain functionality can be delegated to tools and (sub)agents. Tools are deterministic capabilities with a specific bounded function (e.g., run_python(code) or search_web(query)), invoked by the underlying model. We distinguish such fine-grained tools from full-fledged agentic tools (e.g., Claude Code). (Sub)agents are goal-directed components that interpret a user goal, decompose it into substeps, select and execute tools, and iteratively adjust their plan in an “agent loop.” Initially, agentic AI coding tools such as Claude Code and OpenAI Codex implemented one central agent loop steered by a foundation model, followed by conversational tools such as GitHub Copilot and Cursor offering similar capabilities in an “agent mode.” More recently, tool vendors introduced extension and configuration mechanisms that allow developers to write their own (sub)agents that operate in parallel to the central agent loop.
We define a configuration mechanism as a means for developers to tailor tool and agent behavior to a project or workflow (e.g., context files or dedicated subagents). A configuration artifact is a tangible instance of a mechanism (e.g., the actual Markdown file). For example, files like AGENTS.md or CLAUDE.md act as “READMEs for agents” with context-specific information about build commands, coding conventions, and rules for CI/CD pipelines (Mohsenimofidi et al., 2026; OpenAI, 2025). Configuration artifacts are version-controlled, inspectable, and collaboratively maintained. With the increased availability and diversity of agentic tools, the number of configuration mechanisms and related artifacts has also increased.
@@
With this paper, we provide the following contributions: (1) We systematically documented eight configuration mechanisms: Context Files, Skills, Subagents, Commands, Rules, Settings, Hooks, and MCP servers. We identified these from the documentation of Claude Code, GitHub Copilot, Cursor CLI, Gemini CLI, and Codex CLI. Some mechanisms are available in all tools; others are specific to particular tools. (2) We analyzed the adoption of these configuration mechanisms in 2,923 GitHub repositories. Context Files (Markdown files that provide contextual project information) dominated and are often the sole configuration mechanism. Claude Code users apply the broadest range of configuration mechanisms. (3) We analyzed the adoption of Context Files, Skills (bundles of information and tools), and Subagents (specialized agents with their own model context), in more detail. Three Context File formats showed the most dynamic development, with AGENTS.md emerging as a standard. Although Skills and Subagents can be defined for a wide range of purposes, most repositories that adopt them define only one or two artifacts. Moreover, Skills primarily rely on static resources rather than executable scripts to extend tool and agent behavior.
@@ 2. Related Work
Prior work has examined how repository-level context files, prompts, and structured context mechanisms shape agent behavior, efficiency, and integration into development workflows, but has typically focused on single artifact types in isolation. There has been no systematic examination of the broader configuration mechanisms to customize agentic behavior. Independent of tool architecture, context engineering allows developers to design, structure, and supply task- and project-specific information to guide the behavior of tools (Mei et al., 2025; Schmid, 2025; Horthy, 2025; Sahoo et al., 2024; DAIR.AI Prompt Engineering Guide, 2025). Empirical work has mainly focused on repository-maintained context files, such as AGENTS.md. Mohsenimofidi et al. (Mohsenimofidi et al., 2026) analyzed the emergence of context engineering practices in open-source projects and showed that repository-level context files serve as persistent configuration mechanisms that encode architectural constraints, build commands, and workflow conventions. Chatlatanagulchai et al. (Chatlatanagulchai et al., 2025) found that context files are actively maintained through frequent, small changes and focus predominantly on development instructions such as build and testing procedures. Santos et al. (Santos et al., 2025) analyzed 328 CLAUDE.md files and found that architectural concerns dominate, often co-occurring with development guidelines, testing, and project overview.
@@
Villamizar et al. (Villamizar et al., 2025) argued that prompts should be treated as software engineering artifacts and outlined a research agenda for their evolution, reuse, and governance. Zhang et al. (Zhang et al., 2025) introduced a framework that treats contexts as evolving instructions and addresses brevity bias and context collapse via incremental updates.
Recent work further generalized this perspective by formalizing context as a programmable abstraction. Zhang and Wang (Zhang and Wang, 2025) introduced Monadic Context Engineering, modeling context construction and transformation through monadic composition. McMillan (McMillan, 2026) evaluated structured file-native context schemas and showed how format and multi-file organization influence retrieval accuracy and downstream agent performance. Ye et al. (Ye et al., 2026) proposed Meta Context Engineering, a framework in which agentic skills and context artifacts co-evolve via evolutionary search. Jiang and Nam (Jiang and Nam, 2026) explored Cursor rules as one particular configuration mechanism and developed a taxonomy of project contexts.
However, a cross-tool perspective mapping configuration mechanisms and studying their adoption patterns across repositories is missing. Our work addresses this gap.
@@ 3. Data Collection and Analysis
We collected OSS projects hosted on GitHub, selecting repositories belonging to “engineered” software projects (Munaiah et al., 2017) using an updated version of a selection approach from prior work, starting with the SEART GitHub search tool (Mohsenimofidi et al., 2026; SEART, 2025; Dabic et al., 2021). To ensure sufficient project maturity and development history, we selected non-fork repositories with at least two contributors and a license that were created before 1 January 2024 with commits since 1 June 2025. We then excluded archived, disabled, or locked repositories. We applied a licensing filter and a language filter based on popularity, selecting Python, TypeScript, JavaScript, Go, Java, C++, Rust, PHP, C#, and C, and applied filters based on commit and watcher count based on prior research (Mohsenimofidi et al., 2026). This resulted in a sample of 37,249 repositories.
@@
a Copilot also supports CLAUDE.md, AGENTS.md, and GEMINI.md; b Configured via the GitHub web UI, not via files in the project repository;
@@
Table 2. Release dates of agentic AI coding tools and repositories ($n=2,923$; one repository can use multiple tools).
@@
Claude Code 02/2025 (CLI & agents since release) 1,305
@@
957
AGENTS.mdd 05/2025 (Adoption in Codex)
08/2025 (Specification (agentsmd community, 2025))
528
@@
178
Codex CLI 04/2025 (CLI & agents since release) 4
d Repositories using AGENTS.md without any tool-specific configuration artifact.
Table 3. Repository metadata by agentic tool. Cells show median, IQR, and Cliff’s $\delta$ for tool-adopting repositories, each compared against the complement of non-adopting repositories. Significant differences (BH-adjusted) are in bold.
@@
All 6.7 42 2,108 39k
(nn=2,923) (4.3–9.4) (19–105) (965–5,065) (9,501–132k)
Claude 6.2∗∗∗ 40 2,217 42k
(nn=1,305) (4.1–9.2) (19–111) (989–5,270) (9,979–152k)
$\delta$=-0.08 $\delta$=-0.00 $\delta$=0.03 $\delta$=0.04
Copilot 7.1∗∗∗ 42 2,231 45k
(nn=957) (5.0–9.7) (19–113) (1,001–5,607) (9,958–143k)
$\delta$=0.11 $\delta$=0.03 $\delta$=0.04 $\delta$=0.04
AGENTS.md 7.1 44 1,809∗∗∗ 31k∗∗∗
(nn=528) (4.4–9.4) (19–89) (869–3,972) (8,511–91k)
$\delta$=0.03 $\delta$=-0.02 $\delta$=-0.10 $\delta$=-0.10
Cursor 5.5∗∗∗ 52 2,780∗∗∗ 75k∗∗∗
(nn=327) (3.4–7.9) (23–118) (1,264–6,029) (21k–198k)
$\delta$=-0.20 $\delta$=0.09 $\delta$=0.14 $\delta$=0.22
Gemini 6.8 63∗∗∗ 3,302∗∗∗ 58k
(nn=178) (4.7–9.5) (31–142) (1,282–9,711) (15k–189k)
$\delta$=0.02 $\delta$=0.20 $\delta$=0.20 $\delta$=0.12
Codex 7.6 46 6,014 505k
(nn=4) (4.6–11) (32–62) (4,997–6,514) (153k–939k)
$\delta$=0.12 $\delta$=-0.07 $\delta$=0.48 $\delta$=0.66
Note: Mann–Whitney U test (two-sided) comparing each tool’s adopters against non-adopters (complement), with Benjamini–Hochberg FDR correction (6 categories $\times$ 4 metrics = 24 comparisons). Effect sizes: Cliff’s delta. $p<.05$; ∗∗ $p<.01$; ∗∗∗ $p<.001$ (adjusted). Interpret Codex results (nn=4) with caution due to sample size.
Figure 2. Adoption of agentic tools per programming language; AGENTS.md denotes repositories using only that file with no tool-specific configuration. Percentages relative to repository count per tool. Column totals can exceed the overall repository count (multiple tools per repository).
@@
Figure 3. Usage of configuration mechanisms across agentic tools. Repositories using only AGENTS.md without tool-specific configuration are excluded. Percentages relative to repository count per tool. Column totals can exceed the overall repository count (multiple tools per repository).
Diagram showing usage of configuration mechanisms across agentic tools.
Figure 1 outlines the data collection pipeline. We cloned the repositories and searched their default branch for a README file, excluding 11 that became unavailable and 60 without a README. We then used the lingua-language-detector Python library to detect each file’s language, excluding 994 repositories with non-English README files.
For the remaining 36,184 repositories, we used a GPT-5.2-based classification pipeline to determine, based on README content, whether each project qualifies as “engineered” per Munaiah et al. (Munaiah et al., 2017), that is, showing clear evidence of software engineering practices (e.g., a stated purpose, installation and deployment instructions, documentation, testing, CI/CD) and maintenance. The supplementary material includes the prompt (iteratively developed and tested on subsets) and model configuration (OpenAI defaults) (Authors, 2026). The pipeline labeled 32,564 repositories as engineered and excluded the remaining 3,620 (1,416 classified as non-engineered, 2,204 “unsure”). We spot-checked randomly selected samples from each category during prompt development and for the final sample. We then cloned the remaining 32,564 repositories and applied heuristics based on file names and file paths to detect the usage of AI coding tools and configuration mechanisms (see Table 1). These heuristics are based on the answers to RQ1 and are briefly discussed in Section 4. This resulted in 2,923 repositories that use one or more AI coding tools (three additional repositories contained only empty configuration artifacts and were excluded). Our data collection and analysis scripts and the analyzed data are available online (Authors, 2026).
@@ 4. Configuration Mechanisms (RQ1)
We selected five agentic AI coding tools (Table 2) based on the 2025 Stack Overflow Developer Survey (Stack Exchange Inc., 2026): the four most popular AI tools among the surveyed developers, substituting Codex for ChatGPT as the agentic tool by the same vendor, plus Cursor, which has recently been studied in software engineering (He et al., 2026; Jiang and Nam, 2026). Our study focuses on the CLI-based agentic interfaces that first appeared in 2025; for some tools, prior non-agentic versions exist. One author systematically reviewed each tool’s online documentation, documenting configuration mechanisms together with repository-level files and directories that indicate their usage. This was then cross-checked by two other authors and discussed among all three.
We used these findings to develop heuristics for detecting tool usage and configuration mechanisms in repositories (Table 1). For example, a .claude directory or a CLAUDE.md file indicates Claude Code usage; a .claude/agents/ directory with Markdown files indicates Subagents. AGENTS.md is a special case: as a tool-agnostic standard, it is supported by multiple tools, so repositories using only AGENTS.md without tool-specific artifacts are tracked separately (see Table 2, footnote d). Note that for Copilot, Cursor, and Gemini, certain detected files apply to both conversational and agentic interfaces. All documents and the Python scripts that implement our matching strategy are part of the supplementary material (Authors, 2026).
@@
RQ1 (Summary):
  • We identified eight configuration mechanisms spanning a spectrum from static context (e.g., Context Files) to executable and external integrations (e.g., Skills, MCP). Two mechanisms (Context Files and Skills) are supported by all five tools.
  • Despite this convergence, no single tool implements all eight mechanisms.